OpenStudy (anonymous):

A tank contains 1000 liters (L) of a solution consisting of 100 kg of salt dissolved in water. Pure water is pumped into the tank at the rate of 5 L/s, and the mixture - kept uniform by stirring - is pumped out at the same rate. How long will it be until only 10 kg of salt remains in the tank? I need help with this problem plz.

OpenStudy (anonymous):

\[\begin{align*}\frac{dA(t)}{dt}&=\text{(rate in)}-\text{(rate out)}\\ &=\left(0\frac{\text{kg}}{\text{L}}\right)\left(5\frac{\text{L}}{\text{s}}\right)-\left(\frac{A(t)}{1000}\frac{\text{kg}}{\text{L}}\right)\left(5\frac{\text{L}}{\text{s}}\right)\\\\ &=-\frac{1}{200}A(t) \end{align*}\] Do you see how we get this setup? The equation is separable, so finding the solution should be fairly straightforward.

OpenStudy (anonymous):

could you plz finish it so i can get the idea

OpenStudy (anonymous):

Are you telling me you don't know how to solve this equation: \[y'=-\frac{1}{200}y~~?\] It's separable, consider the differential notation for the derivative: \[\begin{align*} \frac{dy}{dx}&=-\frac{1}{200}y\\\\ \frac{dy}{y}&=-\frac{dx}{200}\\\\ \int\frac{dy}{y}&=-\frac{1}{200}\int dx \end{align*}\]

OpenStudy (anonymous):

so i will get y=Ce^(-x/200) and C=100 y(x)=100 e^(-x/200)

OpenStudy (anonymous):

Correct